OpenStudy (anonymous):

Solve -4(x + 1) - 3 = -3(x - 4).

OpenStudy (anonymous):

Do you know how to distribute?

OpenStudy (anonymous):

yes

Mehek (mehek14):

use distributive property \(-4x-4-3=-3x+12\) next combine the x's \(-4-3=x+12\) \(-7=x+12\) subtract 12 \(-7-12=-19\) \(x=-19\)

OpenStudy (anonymous):

do I have to distribute -4 with -3 as well

OpenStudy (anonymous):

No, you do not.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

Zoya has to earn at least $300 to meet her fundraising goal. She has only 100 bracelets that she plans to sell at $5 each. Which inequality shows the number of bracelets, x, Zoya can sell to meet her goal?

OpenStudy (anonymous):